40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 9 12:35:47 2017, gautam, Update, ALS, Fiber ALS IMG_3942.JPG
    Reply  Mon Aug 14 16:24:09 2017, gautam, Update, ALS, Fiber ALS IMG_7471.JPGIMG_7472.JPGIMG_7473.JPGIMG_7474.JPG
       Reply  Wed Aug 16 20:24:23 2017, gautam, Update, ALS, Fiber ALS IMG_7476.JPGIMG_7477.JPGIMG_7478.JPGIMG_7479.JPG
          Reply  Wed Aug 23 17:22:36 2017, gautam, Update, ALS, Fiber ALS - reinstalled 
             Reply  Fri Aug 25 15:54:14 2017, gautam, Update, ALS, Fiber ALS noise measurement IMG_7497.JPGFOL_schematic.pdf20170825_IR_ALS.pdf
                Reply  Fri Aug 25 17:11:07 2017, rana, Update, ALS, Fiber ALS noise measurement 
                   Reply  Tue Sep 26 19:10:13 2017, gautam, Update, ALS, Fiber ALS setup neatened IMG_7605.JPG
                Reply  Sun Aug 27 11:57:31 2017, rana, Update, ALS, Fiber ALS noise measurement 
                   Reply  Tue Aug 29 02:08:39 2017, gautam, Update, ALS, Fiber ALS noise measurement 
                      Reply  Fri Sep 1 19:15:40 2017, gautam, Update, ALS, Fiber ALS noise measurement IR_ALS_noise.pdf
Message ID: 13222     Entry time: Wed Aug 16 20:24:23 2017     In reply to: 13204     Reply to this: 13246
Author: gautam 
Type: Update 
Category: ALS 
Subject: Fiber ALS 

Today, with Johannes' help, I cleaned the fiber tips of the photodiodes. The effect of the cleaning was dramatic - see Attachments #1-4, which are X Beat PD, axial illumination, X Beat PD, oblique illumination, Y beat PD, axial illumination, Y beat PD, oblique illumination. They look much cleaner now, and the feature that looked like a scratch has vanished.

The cleaning procedure followed was:

  • Blow clean air over the fiber tip
  • First, we tried cleaning with the Q-tip like tool, but the results weren't great. The way to use it is to dip the tip in the cleaning solvent for a few seconds, hold the tip to the fiber taking into account the angled cut, and apply 10 gentle quarter turns.
  • Next, we tried cleaning with the wipes. We peeled out an approximately 5" section of the wipe, and laid it out on the table. We then applied cleaning solvent liberally on the central area where we were sure we hadn't touched the wipe. Then you just drag the fiber tip along the soaked part of the wipe. If you get the angle exactly right, the fiber glides smoothly along the surface, but if you are a little misaligned, you get a scratchy sensation. 
  • Blow dry and inspect.

I will repeat this procedure for all fiber connections once I start putting the box back together - I'm almost done with the new box, just waiting on some hardware to arrive.

 

Quote:

Today, I borrowed the fiber microscope from Johannes and took a look at the fibers coupled to the PDs. The PD labelled "BEAT PD AUX Y" has an end that seems scratched (Attachments #1 and #2). The scratch seems to be on (or at least very close to) the core. The other PD (Attachments #3 and #4) doesn't look very clean either, but at least the area near the core seems undamaged. The two attachments for each PD corresponds to the two available lighting settings on the fiber microscope.

I have not attempted to clean them yet, though I have also borrowed the cleaning supplies to facilitate this from Johannes. I also plan to inspect the ends of all other fiber connections before re-installing them.

 

Attachment 1: IMG_7476.JPG  1.265 MB  | Hide | Hide all | Show all
IMG_7476.JPG
Attachment 2: IMG_7477.JPG  1.307 MB  | Hide | Hide all | Show all
IMG_7477.JPG
Attachment 3: IMG_7478.JPG  1.296 MB  | Hide | Hide all | Show all
IMG_7478.JPG
Attachment 4: IMG_7479.JPG  1.247 MB  | Show | Hide all | Show all
ELOG V3.1.3-